013 (part 1 of 4) 10.0 points You throw a softball (of mass 200 g) straight up into the air. It reaches a maximum altitude of 13.9 m and then returns to you What is the gravitational potential energy of the soft ball at its highest position? Assume the ball departed from and returned to ground level. Answer in units of J.
014 (part 2 of 4) 10.0 points Assume no energy is lost by the softball while it is in the air What is the kinetic energy of the soft ball as soon as it leaves your hand? Answer in units of J 015 (part 3 of 4) 10.0 points What is the kinetic energy of the soft ball when it returns to your hand? Answer in units of J. 016 (part 4 of 4) 10.0 points What is the speed of the ball? Answer in units of m/s.
Answer #1

1) Gravitational potential energy of mass m at height, h =mgh

Hence gravitational potential energy of ball at highest position

= 0.2*9.8*13.9 = 27.2 J

2) As no energy is lost in air, kinetic energy of ball just after leaving the hand is converted to potential energy at highest point. Hence kinetic energy of ball just after leaving the hand

= 27.2 J

3) Kinetic energy of ball as it returns, is again 27.2 J, as there is no change in it's potential energy on returning and there is no loss of energy in air.

4) Kinetic energy = mv^2/2

27.2 = 0.2*v^2/2

v = 16.5 m/s